OpenStudy (anonymous):

Evaluate 3^2 + (7 - 2) • 4 - 6 over 3. 24 54 27 28

OpenStudy (anonymous):

i love ur profile pic! #JESUSFORLIFE!!!!

OpenStudy (anonymous):

Haha Thanks @WhiteFangIII

OpenStudy (igreen):

\(3^2 + (7 - 2) \cdot 4 - 6\) Okay, we have to follow PEMDAS. Parenthesis. Exponent. Multiplication.---- | --------------- Whichever comes first from left to right. Division.--------- Addition. --------- | ------------- Whichever comes first from left to right. Subtraction. ------- So our first step would be the Parenthesis, or to subtract 7 -2.

OpenStudy (igreen):

* 7 - 2

OpenStudy (anonymous):

np

OpenStudy (anonymous):

5

OpenStudy (igreen):

Yep, so that gives us: \(3^2 + 5 \cdot 4 - 6\) Now we do Exponent, so what's \(3^2\)? @shybearr15

OpenStudy (anonymous):

9

OpenStudy (igreen):

Yep, that gives us: \(9 + 5 \cdot 4 - 6\) Now multiply 5 * 4.

OpenStudy (igreen):

@shybearr15

OpenStudy (anonymous):

20

OpenStudy (igreen):

Yep, that gives us: \(9 + 20 - 6\) Now add 9 + 20, then subtract 6.

OpenStudy (anonymous):

24. Yay I got it right in the first place haha. Thank you.

OpenStudy (igreen):

No problem.

OpenStudy (igreen):

Wait..no you made a mistake..

OpenStudy (igreen):

Add 20 + 9..

OpenStudy (igreen):

20 + 9 = 29 29 - 6 = 23..but that's not one of your options..

OpenStudy (anonymous):

oh... then... whaaaat

OpenStudy (igreen):

Oh, wait, it says over 3.. so is it \(\dfrac{3^2 + (7 - 2) \cdot 4 - 6}{3}\) or is it \(3^2 + (7 - 2) \cdot 4 - \dfrac{6}{3}\) @shybearr15

OpenStudy (igreen):

Oh, I'm pretty sure it's 6 / 3.. 6 / 3 = 2. So what's 29 - 2? @shybearr15

OpenStudy (igreen):

@shybearr15

OpenStudy (anonymous):

27 Sorry my computer being stupid and shut off for while